A plane flies 285 km due west and then 320 km due north. what are its distance and bearing from the starting point?
thanks

Answer by fractalier(1829)

The distance is found via the Pythagorean Theorem...
c^2 = a^2 + b^2
c^2 = 285^2 + 320^2
so that
c = 428.5
The angle is arc tan(320/285) = 48.3° so that bearing is 180 - 48.3 =
131.7°